Hello, I would like to receive help with following question,
I hope you can help!
Maria

Zhra Manufacturing Company has two​ departments, machining and finishing. For a given​ period, the following costs were incurred by the company as a​ whole: direct​material, £350,000. direct​ labour, £105,000 and indirect​ production, £40,000. The grand total was £495,000. The machining department incurred 70% of the​direct-material costs, but only 20% of the​ direct-labour costs. As is​ commonplace, indirect production costs incurred by each department were allocated to products in proportion to the​ direct-labour costs of products within the departments. Three products were produced.

Product

Direct Material

Direct Labour

Sigma

30%

30%

Chi

10%

10%

Delta

60%

60%

Total for machining department

100%

100%

Sigma

30%

20%

Chi

30%

30%

Delta

40%

50%

Total for finishing department

100%

100%

The indirect production costs incurred by the machining and finishing departments and allocated to all products therein amounted to​ machining, £22,000 and​ finishing, £18,000.

Requirement 1. Compute the total costs incurred by the machining department and added by the finishing department.

Machining

Finishing

Total

Direct materials

Direct labour

Indirect production

Total


Thank you!
Answer #1
Machining Finishing Total
Direct materials 245000 105000 350000
Direct labour 21000 84000 105000
Indirect production 22000 18000 40000
Total 288000 207000 495000
Formulas used:
Machining Finishing
Direct materials =350000*70% =350000*30%
Direct labour =105000*20% =105000*80%